If you toss a fair coin and the probability of heads is 1/2, what is the probability of NOT getting heads?
A0
B1/4
C1/2
D1
Answer & Solution
Correct answer: C. 1/2
**For any event A:** P(A) + P(not A) = 1 (the sum of all probabilities equals 1).
**So P(not heads) = 1 − 1/2 = 1/2. (Equivalently, P(tails) = 1/2 — same thing for a 2-outcome experiment.) This is the 'complement rule'.**
Keep practising Algebra: Logarithms
Browse all Algebra: Logarithms questions for GMAT, or explore more Mathematics topics.
Related questions
Probability bingoSpot the odd one outPick the probabilityProbability is measured on a scale from:Most everyday events have a probability that is:Experimental and theoretical probability differ most when trials are:A probability of 0.75 means a chance of:A spinner resting on 1 to 8 gives outcomes that are: